OpenStudy (anonymous):

At the start of an experiment, there are 50 bacteria in a dish. The bacteria population is expected to grow at a rate of 220% each day. Which is the best prediction for the baceria population after 8 days?

OpenStudy (calculusxy):

Convert 220% into decimal. What do you get?

OpenStudy (anonymous):

2.2

OpenStudy (calculusxy):

Okay. So we can make an equation from it. Such as y = mx + b

OpenStudy (calculusxy):

Do you know what that is?

OpenStudy (anonymous):

ok got it. thanks

OpenStudy (calculusxy):

What's your answer?

OpenStudy (calculusxy):

Just to make sure, don't have `y = mx + b` as your answer

OpenStudy (calculusxy):

I mean that would be the format of your answer. Do you understand?

Directrix (directrix):

y = 50 * (1 + 2.2)^8